if 18.8 ml of 1.25 m sodium hydroxide solution is necessary to neutralize 25.0 ml of hydrochloric acid solution to a phenolphthalein end point, calculate the molarity of the hydrochloric acid solution

1
Expert's answer
2022-05-27T12:58:03-0400

HCl + NaOH = NaCl + H2O

1. Using the formula n = CM * V , we find how many moles of NaOH. In the formula,

n — amount of solute (mol);

CM — concentration of molarity (mol/l);

V — volume of solution (litres):


n = 1,25 * 0,0188 = 0,0235 mol NaOH


2. According to the above reaction, 1 mol of NaOH reacts with 1 mol of HCl. We find how many moles of HCl react with 0.0235 mol NaOH:

1mol NaOH — 1mol HCl

0,0235mol NaOH — X mol HCl

X = 0,0235mol HCl


3. We find the molar concentration of HCl using the following formula:

CM = n / V = 0,0235 mol / 0,025 litres = 0,94

ANSWER: 0,94 mol/l HCl